the probability function of a discrete random variable x is given by p(x) = q x x = 1 2 3 4

1. value of q
2. probability distribution table of x
3. mean
1
Expert's answer
2020-05-31T19:17:26-0400
"p(x)=qx, x=1,2,3,4"

1. value of "q"


"q(1)+q(2)+q(3)+q(4)=1""q=0.1"

2. probability distribution table of x 


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 1 & 2 & 3 & 4 \\ \hline
 p(x) & 0.1 & 0.2 & 0.3 & 0.4 \\
 
 
\end{array}"

3. mean


"mean=\mu_X=E(X)=""=1(0.1)+2(0.2)+3(0.3)+4(0.4)=3"